LaTanya Corporation is planning to issue bonds with a face value of $101,500 and a coupon rate of 8 percent. The bonds mature in seven years. Interest is paid annually on December 31. All of the bonds will be sold on January 1 of this year. (FV of $1, PV of $1, FVA of $1, and PVA of $1) (Use the appropriate factor(s) from the tables provided. Round your final answer to whole dollars.)

Required: Compute the issue (sale) price on January 1 of this year for each of the following independent cases:

a. Case A: Market interest rate (annual): 8 percent.

issue price=

b. Case B: Market interest rate (annual): 6 percent.

issue price=

c. Case C: Market interest rate (annual): 9 percent.

issue price=
Answer #1

Formulas to be noted

PV of the future cash flow = future cash flow / (1+r)^n

PV of annuity = periodic payment × {(1-(1+r)^-n)/r}

Issue price of bond = PV of future cash flow + PV of annuity

r = interest rate ; n = years to maturity

Case A

Market interest rate = 8% ,

Interest amount = 101500*0.08 = 8120

Issue price = (101500 * 0.58349) + (8120 * 5.20637)

=59224.28 + 42275.72

=$ 101500

CASE B

MARKET INTEREST =6%

ISSUE PRICE = (101500* 0.665057) + ( 8120*5.582381)

= 67503.3 + 45328.94

=$112832.2

CASE C

Market interest rate = 9%

Issue price = (101500* 0.547034) + (8120* 5.032953)

=55523.98 + 40867.58

= $ 96391.55